### v1

# Online Posterior Sampling with a Diffusion Prior

Abstract
Posterior sampling in contextual bandits with a Gaussian prior can be implemented exactly or approximately using the Laplace approximation. The Gaussian prior is computationally efficient but it cannot describe complex distributions. In this work, we propose approximate posterior sampling algorithms for contextual bandits with a diffusion model prior. The key idea is to sample from a chain of approximate conditional posteriors, one for each stage of the reverse process, which are estimated in a closed form using the Laplace approximation. Our approximations are motivated by posterior sampling with a Gaussian prior, and inherit its simplicity and efficiency. They are asymptotically consistent and perform well empirically on a variety of contextual bandit problems.
1 Introduction
A multi-armed bandit [26, 6, 29] is an online learning problem where an agent sequentially interacts with an environment over $n$ rounds with the goal of maximizing its rewards. In each round, it takes an action and receives its stochastic reward. The mean rewards of the actions are unknown a priori and must be learned. This leads to the exploration-exploitation dilemma: explore actions to learn about them or exploit the action with the highest estimated reward. Bandits have been successfully applied to problems where uncertainty modeling and adaptation to it are beneficial, such as in recommender systems [31, 53, 24, 34] and hyper-parameter optimization [33].
Contextual bandits [28, 31] with linear [13, 1] and generalized linear models (GLMs) [16, 32, 2, 25] have become popular due to the their flexibility and efficiency. The features in these models can be hand-crafted or learned from historic data [39], and the models can be also updated incrementally [1, 23]. While the original algorithms for linear and GLM bandits were based on upper confidence bounds (UCBs) [13, 1, 16], Thompson sampling (TS) is more popular in practice [11, 3, 41, 43]. The key idea in TS is to explore by sampling from the posterior distribution of model parameter $\theta_{*}$. TS uses the prior knowledge about $\theta_{*}$ to speed up exploration [11, 39, 35, 9, 20, 19, 5]. When the prior is a multivariate Gaussian, the posterior of $\theta_{*}$ can be updated and sampled from efficiently [11]. Unfortunately, this prior has limited expressive power, because it cannot even represent multimodal distributions. To address this limitation, we study posterior sampling with complex priors.
We make the following contributions. First, we propose novel posterior sampling approximations for linear models and GLMs with a diffusion model prior. The key idea is to sample from a chain of approximate conditional posteriors, one for each stage of the reverse process, which are estimated in a closed form. In linear models, each conditional is a product of two Gaussians, representing prior knowledge and diffused evidence (Theorem 2). In GLMs, each conditional is obtained by a Laplace approximation, which mixes prior knowledge and evidence (Theorem 4). Our approximations are motivated by posterior sampling with a Gaussian prior, and inherit its simplicity and efficiency. The prior works (Section 7) rely on the score of the likelihood of observations. In online learning, the likelihood score tends to infinity as the number of observations increases, which causes divergence of the existing approximations (Section 6.2).
Our second major contribution is in theory. More specifically, we prove that our posterior approximation is asymptotically consistent (Theorem 3). The key idea in the proof is that the conditional posteriors in Theorem 2 concentrate at a scaled unknown model parameter as the number of observations increases. While this claim is asymptotic, it is an expected property of a posterior distribution. Many prior works, such as Chung et al. [12], do not poses this property.
Our last contribution is empirical evaluation on contextual bandit problems. We focus on bandits because the ability to represent all levels of uncertainty well is critical for exploration. Our experiments (Section 6.2) show that a score-based method fails to do so. Note that our posterior approximations are general and not restricted to bandits.
2 Setting
We start with introducing our notation. Random variables are capitalized, except for Greek letters like $\theta$. We denote the marginal and conditional probabilities under probability measure $p$ by $p(X=x)$ and $p(X=x\mid Y=y)$, respectively. When the random variables are clear from context, we write $p(x)$ and $p(x\mid y)$. We denote by $X_{n:m}$ and $x_{n:m}$ a collection of random variables and their values, respectively. For a positive integer $n$, we define $[n]=\left\{1,\dots,n\right\}$. The indicator function is $\mathds{1}\!\left\{\cdot\right\}$. The $i$-th entry of vector $v$ is $v_{i}$. If the vector is already indexed, such as $v_{j}$, we write $v_{j,i}$. We denote the maximum and minimum eigenvalues of matrix $M\in\mathbb{R}^{d\times d}$ by $\lambda_{1}(M)$ and $\lambda_{d}(M)$, respectively.
The posterior sampling problem can be formalized as follows. Let $\theta_{*}\in\Theta$ be an unknown model parameter and $\Theta\subseteq\mathbb{R}^{d}$ be the space of model parameters. Let $h=\left\{(\phi_{\ell},y_{\ell})\right\}_{\ell\in[N]}$ be the history of $N$ noisy observations of $\theta_{*}$, where $\phi_{\ell}\in\mathbb{R}^{d}$ is the feature vector for $y_{\ell}\in\mathbb{R}$. We assume that
$y_{\ell}=g(\phi_{\ell}^{\top}\theta_{*})+\varepsilon_{\ell}\,,$ (1)
where $g:\mathbb{R}\to\mathbb{R}$ is the mean function and $\varepsilon_{\ell}$ is independent zero-mean $\sigma^{2}$-sub-Gaussian noise for $\sigma>0$. Let $p(h\mid\theta_{*})$ be the likelihood of observations in history $h$ under model parameter $\theta_{*}$ and $p(\theta_{*})$ be its prior probability. By Bayes’ rule, the posterior distribution of $\theta_{*}$ given $h$ is
$p(\theta_{*}\mid h)\propto p(h\mid\theta_{*})\,p(\theta_{*}).$ (2)
We want to sample from $p(\cdot\mid h)$ efficiently when the prior distribution is represented by a diffusion model. As a stepping stone, we review existing posterior formulas for multivariate Gaussian priors. This motivates our solution for diffusion model priors.
2.1 Linear Model
The posterior of $\theta_{*}$ in linear models has the following form. Let $p(\theta_{*})=\mathcal{N}(\theta_{*};\theta_{0},\Sigma_{0})$ be the prior distribution, where $\theta_{0}\in\mathbb{R}^{d}$ and $\Sigma_{0}\in\mathbb{R}^{d\times d}$ are the prior mean and covariance, respectively.
Assumption 1.
Let $g$ in (1) be an identity and $\varepsilon_{\ell}\sim\mathcal{N}(0,\sigma^{2})$. Then the likelihood of $h$ under model parameter $\theta_{*}$ is $p(h\mid\theta_{*})\propto\exp[-\sigma^{-2}\sum_{\ell=1}^{N}(y_{\ell}-\phi_{\ell}^{\top}\theta_{*})^{2}]$.
Under these assumptions, $p(\theta_{*}\mid h)\propto\mathcal{N}(\theta_{*};\hat{\theta},\hat{\Sigma})$ [10], where
$\hat{\theta}=\hat{\Sigma}(\Sigma_{0}^{-1}\theta_{0}+\bar{\Sigma}^{-1}\bar{\theta})\,,\quad\hat{\Sigma}=(\Sigma_{0}^{-1}+\bar{\Sigma}^{-1})^{-1}\,,$ (3)
$\bar{\theta}=\sigma^{-2}\bar{\Sigma}\sum_{\ell=1}^{N}\phi_{\ell}y_{\ell}$, and $\bar{\Sigma}^{-1}=\sigma^{-2}\sum_{\ell=1}^{N}\phi_{\ell}\phi_{\ell}^{\top}$. Therefore, the posterior of $\theta_{*}$ is a product of two multivariate Gaussians, representing the prior knowledge about $\theta_{*}$, given by $\mathcal{N}(\theta_{0},\Sigma_{0})$; and empirical evidence, given by $\mathcal{N}(\bar{\theta},\bar{\Sigma})$.
2.2 Generalized Linear Model
Algorithm 1 $\tt IRLS$: Iteratively reweighted least squares.
1:Input: Prior parameters $\theta_{0}$ and $\Sigma_{0}$, history of observations $h=\left\{(\phi_{\ell},y_{\ell})\right\}_{\ell\in[N]}$
2:
3:Initialize $\hat{\theta}\in\mathbb{R}^{d}$
4:repeat
5:     for stage $\ell=1,\dots,N$ do $z_{\ell}\leftarrow\phi_{\ell}^{\top}\hat{\theta}+(y_{\ell}-g(\phi_{\ell}^{\top}\hat{\theta}))/\dot{g}(\phi_{\ell}^{\top}\hat{\theta})$
6:     $\hat{\Sigma}\leftarrow\left(\Sigma_{0}^{-1}+\sum_{\ell=1}^{N}\dot{g}(\phi_{\ell}^{\top}\hat{\theta})\phi_{\ell}\phi_{\ell}^{\top}\right)^{-1}\,,\ \hat{\theta}\leftarrow\hat{\Sigma}\left(\Sigma_{0}^{-1}\theta_{0}+\sum_{\ell=1}^{N}\dot{g}(\phi_{\ell}^{\top}\hat{\theta})\phi_{\ell}z_{\ell}\right)$
7:until $\hat{\theta}$ converges
8:
9:Output: Posterior mean $\hat{\theta}$ and covariance $\hat{\Sigma}$
Generalized linear models (GLMs) [36] extend linear models (Section 2.1) to non-linear monotone mean functions $g$ in (1). For instance, in logistic regression, $g(u)=1/(1+\exp[-u])$ is a sigmoid. The likelihood of observations in GLMs has the following form [25].
Assumption 2.
Let $h=\left\{(\phi_{\ell},y_{\ell})\right\}_{\ell\in[N]}$ be a history of $N$ observations under mean function $g$ and the corresponding noise. Then $\log p(h\mid\theta_{*})\propto\sum_{\ell=1}^{N}y_{\ell}\phi_{\ell}^{\top}\theta_{*}-b(\phi_{\ell}^{\top}\theta_{*})+c(y_{\ell})$, where $c$ is a real function, and $b$ is a function whose derivative is the mean function, $\dot{b}=g$.
The posterior distribution of $\theta_{*}$ in GLMs does not have a closed form in general [10]. Therefore, it is often approximated by the Laplace approximation. Let the prior distribution of the model parameter be $p(\theta_{*})=\mathcal{N}(\theta_{*};\theta_{0},\Sigma_{0})$, as in Section 2.1. Then the Laplace approximation is $\mathcal{N}(\hat{\theta},\hat{\Sigma})$, where $\hat{\theta}$ is the maximum a posteriori (MAP) estimate of $\theta_{*}$ and $\hat{\Sigma}$ is the corresponding covariance.
The MAP estimate $\hat{\theta}$ can be obtained by iteratively reweighted least squares (IRLS) [51], which we present in Algorithm 1. $\tt IRLS$ is an iterative method where the MAP estimate $\hat{\theta}$ is updated (line 5) until convergence. This is guaranteed under mild conditions due to the convexity of the problem. In the updates, $\dot{g}$ is the derivative of the mean function and the pseudo-observation $z_{\ell}$ (line 4) plays the role of the observation $y_{\ell}$ in the empirical mean $\bar{\theta}$. The IRLS solution has a similar structure to (3). Specifically, $\mathcal{N}(\hat{\theta},\hat{\Sigma})$ is a product of two multivariate Gaussians, representing the prior knowledge about $\theta_{*}$ and empirical evidence.
2.3 Towards Diffusion Model Priors
The assumption that $p(\theta_{*})=\mathcal{N}(\theta_{*};\theta_{0},\Sigma_{0})$ is limiting, for instance because it precludes multimodal priors. We relax it by representing $p(\theta_{*})$ by a diffusion model, which we call a diffusion model prior. We propose an efficient posterior sampling approximation for this prior where the prior and empirical evidence are mixed similarly to (3) and $\tt IRLS$ (Section 4). Before we do this, we review diffusion models in Section 3.
3 Diffusion Models
A diffusion model [45, 18] is a generative model trained by diffusing samples from an unknown distribution. While multiple types of diffusion models exist [48], we adopt the probabilistic formulation and presentation of Ho et al. [18]. A diffusion model is a graphical model with $T$ stages indexed by $t\in[T]$. Each stage $t$ is associated with a latent variable $S_{t}\in\mathbb{R}^{d}$. A sample from the model is represented by an observed variable $S_{0}\in\mathbb{R}^{d}$. We visualize the diffusion model in Figure 1. In the forward process, the sample $s_{0}$ is diffused through a sequence of variables $S_{1},\dots,S_{T}$. This process is used to learn the reverse process, where the sample $s_{0}$ is generated through a sequence of variables $S_{T},\dots,S_{0}$. To sample $s_{0}$ from the posterior (Section 4), we add a random variable $H$, which represents partial information about $s_{0}$. We introduce forward and reverse diffusion processes next. Learning of the reverse process is described in Appendix B. While this is a critical component of diffusion models, it is not necessary to introduce our posterior approximations.
Forward process. The forward process diffuses an observed sample $s_{0}$ through a chain of latent variables $S_{1},\dots S_{T}$ (Figure 1). We denote the probability measure under this process by $q$ and define its joint probability distribution as
$\textstyle q(s_{1:T}\mid s_{0})=\prod_{t=1}^{T}q(s_{t}\mid s_{t-1})\,,\quad\forall t\in[T]:q(s_{t}\mid s_{t-1})$ $=\mathcal{N}(s_{t};\sqrt{\alpha_{t}}s_{t-1},\beta_{t}I_{d})\,,$ (4)
where $q(s_{t}\mid s_{t-1})$ is the conditional density of mapping a less diffused $s_{t-1}$ to a more diffused $s_{t}$. The diffusion rate is set by parameters $\alpha_{t}\in(0,1)$ and $\beta_{t}=1-\alpha_{t}$. We also define $\bar{\alpha}_{t}=\prod_{\ell=1}^{t}\alpha_{\ell}$. The forward process is sampled as follows. First, a clean sample $s_{0}$ is chosen. Then $S_{t}\sim q(\cdot\mid s_{t-1})$ are sampled, starting from $t=1$ to $t=T$.
Reverse process. The reverse process maps a diffused sample $S_{T}$ to the observed sample $S_{0}$ through a chain of variables $S_{T},\dots,S_{0}$ (Figure 1). We denote the probability measure under this process by $p$ and define its joint probability distribution as
$\textstyle p(s_{0:T})$ $=p(s_{T})\prod_{t=1}^{T}p(s_{t-1}\mid s_{t})\,,$ (5)
$p(s_{T})$ $=\mathcal{N}(s_{T};\mathbf{0}_{d},I_{d})\,,\quad\forall t\in[T]:p(s_{t-1}\mid s_{t})=\mathcal{N}(s_{t-1};\mu_{t}(s_{t}),\Sigma_{t})\,,$
where $p(s_{t-1}\mid s_{t})$ represents the conditional density of mapping a more diffused sample $s_{t}$ to a less diffused sample $s_{t-1}$. The function $\mu_{t}$ predicts the mean of $S_{t-1}\mid s_{t}$ and is learned from the data (Appendix B). As in Ho et al. [18], we keep the covariance fixed at $\Sigma_{t}=\tilde{\beta}_{t}I_{d}$, where $\tilde{\beta}_{t}$ is given in (13) in Appendix B. This is also called stable diffusion. We make this assumption only to simplify exposition. All our derivations in Section 4 hold when $\Sigma_{t}$ is learned, for instance as in Bao et al. [8].
The process is called reverse because it is learned by reversing the forward process. The reverse process is sampled as follows. First, a diffused sample $S_{T}\sim p$ is chosen. After that, $S_{t-1}\sim p(\cdot\mid s_{t})$ are sampled, starting from $t=T$ to $t=1$.
Forward process (probability measure $q$) Reverse process (probability measure $p$)
$S_{T}\leftarrow S_{T-1}\leftarrow\dots\leftarrow S_{1}\leftarrow S_{0}\to H$ $S_{T}\to S_{T-1}\to\dots\to S_{1}\to S_{0}\to H$
Figure 1: Graphical models of the forward and reverse processes in the diffusion model. The variable $H$ represents partial information about $S_{0}$.
4 Posterior Sampling
This section is organized as follows. We derive a general posterior distribution for a chain of random variables conditioned on observations in Section 4.1. After that, in Sections 4.2 and 4.4, we specialize it to the observation models in Section 2.
4.1 Chain Model Posterior
Let $h=\left\{(\phi_{\ell},y_{\ell})\right\}_{\ell\in[N]}$ denote a history of $N$ observations (Section 2) and $H$ be the corresponding random variable. In this section, we assume that $h$ is fixed. The Markovian structure of the reverse process in Figure 1 implies that the joint probability distribution conditioned on $h$ factors as
$p(s_{0:T}\mid h)=p(s_{T}\mid h)\prod_{t=1}^{T}p(s_{t-1}\mid s_{t},h)\,.$
Therefore, $p(s_{0:T}\mid h)$ can be sampled from efficiently by first sampling from $p(s_{T}\mid h)$ and then from $T$ conditional distributions $p(s_{t-1}\mid s_{t},h)$. We derive these next.
Lemma 1.
Let $p$ be a probability measure over the reverse process (Figure 1). Then
$p(s_{T}\mid h)$ $\propto\textstyle\int_{s_{0}}p(h\mid s_{0})\,p(s_{0}\mid s_{T})\mathop{}\!\mathrm{d}s_{0}\,p(s_{T})\,,$
$\forall t\in[T]\setminus\left\{1\right\}:p(s_{t-1}\mid s_{t},h)$ $\propto\textstyle\int_{s_{0}}p(h\mid s_{0})\,p(s_{0}\mid s_{t-1})\mathop{}\!\mathrm{d}s_{0}\,p(s_{t-1}\mid s_{t})\,,$
$p(s_{0}\mid s_{1},h)$ $\propto p(h\mid s_{0})\,p(s_{0}\mid s_{1})\,.$
Proof.
The claim is proved in Section A.1. ∎
4.2 Linear Model Posterior
Now we specialize Lemma 1 to the diffusion model prior (Section 3) and linear models (Section 2.1). The prior distribution is the reverse process in (5),
$p(s_{T})=\mathcal{N}(s_{T};\mathbf{0}_{d},I_{d})\,,\quad p(s_{t-1}\mid s_{t})=\mathcal{N}(s_{t-1};\mu_{t}(s_{t}),\Sigma_{t})\,.$
The term $p(h\mid s_{0})$ is the likelihood of observations in Assumption 1. To fully instantiate Lemma 1, we make additional assumptions on the integrals in it.
Theorem 2.
Let $p$ be a probability measure over the reverse process (Figure 1). Let $\bar{\theta}$ and $\bar{\Sigma}^{-1}$ be defined as in (3). Suppose that we approximate
$\textstyle\int_{s_{0}}p(h\mid s_{0})\,p(s_{0}\mid s_{t})\mathop{}\!\mathrm{d}s_{0}\approx p(h\mid s_{t}/\sqrt{\bar{\alpha}_{t}})$ (6)
for all $t\in[T]$. Then $p(s_{T}\mid h)\approx\mathcal{N}(s_{T};\hat{\mu}_{T+1}(h),\hat{\Sigma}_{T+1}(h))$, where
[…]
For $t\in[T]$, we have $p(s_{t-1}\mid s_{t},h)\approx\mathcal{N}(s_{t-1};\hat{\mu}_{t}(s_{t},h),\hat{\Sigma}_{t}(h))$, where
$\hat{\mu}_{t}(s_{t},h)=\hat{\Sigma}_{t}(h)(\Sigma_{t}^{-1}\mu_{t}(s_{t})+\bar{\Sigma}^{-1}\bar{\theta}/\sqrt{\bar{\alpha}_{t-1}})\,,\quad\hat{\Sigma}_{t}(h)=(\Sigma_{t}^{-1}+\bar{\Sigma}^{-1}/\bar{\alpha}_{t-1})^{-1}\,.$ (8)
Proof.
The proof is in Section A.2 and comprises four steps. First, we fix $t$ and apply approximation (6). Second, we rewrite the likelihood as in (3). Third, we reparameterize the likelihood, from a function of $s_{t}/\sqrt{\bar{\alpha}_{t}}$ to a function of $s_{t}$. Finally, we combine the likelihood with the Gaussian prior using Lemma 6 in Section A.5. ∎
Algorithm 2 $\tt LaplaceDPS$: Laplace posterior sampling with a diffusion model prior.
1:Input: Diffusion model parameters $(\mu_{t},\Sigma_{t})_{t\in[T]}$, history of observations $h$
2:
3:Initial sample $S_{T}\sim\mathcal{N}(\hat{\mu}_{T+1}(h),\hat{\Sigma}_{T+1}(h))$
4:for stage $t=T,\dots,1$ do $S_{t-1}\sim\mathcal{N}(\hat{\mu}_{t}(S_{t},h),\hat{\Sigma}_{t}(h))$
5:
6:Output: Posterior sample $S_{0}$
The algorithm that samples from the posterior distribution in Theorem 2 is presented in Algorithm 2. We call it Laplace diffusion posterior sampling ($\tt LaplaceDPS$) because its generalization to GLMs utilizes the Laplace approximation. $\tt LaplaceDPS$ works as follows. First, $S_{T}$ is sampled conditioned on evidence $h$ (line 2) from the distribution in (7). This distribution is a product of two Gaussians, one for the prior $\mathcal{N}(\mathbf{0}_{d},I_{d})$ and one for the diffused evidence up to stage $T$, $\mathcal{N}(\sqrt{\bar{\alpha}_{T}}\bar{\theta},\bar{\alpha}_{T}\bar{\Sigma})$. All other variables, $S_{t-1}$ for $t\in[T]$, are sampled conditioned on $s_{t}$ and evidence $h$ (line 3) from the distribution in (8). This distribution is again a product of two Gaussians, one for the conditional prior $\mathcal{N}(\mu_{t}(s_{t}),\Sigma_{t})$ and one for the diffused evidence up to stage $t-1$, $\mathcal{N}(\sqrt{\bar{\alpha}_{t-1}}\bar{\theta},\bar{\alpha}_{t-1}\bar{\Sigma})$. The final output $S_{0}$ is a sample from the diffusion prior.
When compared to Section 2, the prior and empirical evidence are mixed conditionally in a $T$-stage chain. Thus the computational cost of $\tt LaplaceDPS$ is $T$ times higher than that of posterior sampling with a Gaussian prior (Section 2). We validate this empirically in Section 6.2.
4.3 Key Approximation in Theorem 2
Now we motivate our approximation in (6). Using the forward process, $s_{t}=\sqrt{\bar{\alpha}_{t}}s_{0}+\sqrt{1-\bar{\alpha}_{t}}\tilde{\varepsilon}_{t}$ for any $s_{0}$, where $\tilde{\varepsilon}_{t}\sim\mathcal{N}(\mathbf{0}_{d},I_{d})$ is a standard Gaussian noise. This identity can be rearranged as $s_{0}=(s_{t}-\sqrt{1-\bar{\alpha}_{t}}\tilde{\varepsilon}_{t})/\sqrt{\bar{\alpha}_{t}}$. So $s_{0}$ can be viewed as a random variable with mean $s_{t}/\sqrt{\bar{\alpha}_{t}}$, with which we replace it in (6). This approximation is a simplification of Chung et al. [12] where the score component is neglected.
Our approximation has several notable properties. First, $\sqrt{(1-\bar{\alpha}_{t})/\bar{\alpha}_{t}}\to 0$ as $t\to 1$. Therefore, it becomes more precise in later stages of the reverse process. Second, in the absence of evidence $h$, the approximation vanishes, and all posterior distributions in Theorem 2 reduce to the priors in (5). Finally, as the number of observations increases, sampling from the posterior in Theorem 2 is asymptotically consistent.
Theorem 3.
Fix $\theta_{*}\in\mathbb{R}^{d}$. Let $\tilde{\theta}\leftarrow{\tt LaplaceDPS}((\mu_{t},\Sigma_{t})_{t\in[T]},h)$, where $h=\left\{(\phi_{\ell},y_{\ell})\right\}_{\ell\in[N]}$ is a history of $N$ observations. Suppose that $\lambda_{d}(\bar{\Sigma}^{-1})\to\infty$ as $N\to\infty$, where $\bar{\Sigma}$ is defined in (3). Then $\mathbb{P}\left(\lim_{N\to\infty}\|\tilde{\theta}-\theta_{*}\|_{2}=0\right)=1$.
Proof.
The proof is in Section A.3. The key idea is that the conditional posteriors in (7) and (8) concentrate at a scaled unknown model parameter $\theta_{*}$ as the number of observations increases, which we formalize as $\lambda_{d}(\bar{\Sigma}^{-1})\to\infty$. ∎
The bound in Theorem 3 can be interpreted as follows. The sampled parameter $\tilde{\theta}$ approaches the true unknown parameter $\theta_{*}$ as the number of observations $N$ increases. To guarantee that the posterior shrinks uniformly in all directions, we assume that the number of observations in all directions grows linearly with $N$. This is akin to assuming that $\lambda_{d}(\bar{\Sigma}^{-1})=\Omega(N)$. This lower bound can be attained in linear models by choosing feature vectors according to the D-optimal design [38]. Since our proof is asymptotic, we can neglect some finite-time errors and this simplifies the argument.
4.4 GLM Posterior
The Laplace approximation in GLMs (Section 2.2) naturally generalizes the exact posterior distribution in linear models (Section 2.1). We generalize Theorem 2 to GLMs along the same lines.
Theorem 4.
Let $p$ be a probability measure over the reverse process (Figure 1). Suppose that we apply approximation (6) for all $t\in[T]$. Then the distribution of $S_{T}\mid h$ is
$p(s_{T}\mid h)\approx\mathcal{N}(s_{T};\hat{\mu}_{T+1}(h),\hat{\Sigma}_{T+1}(h))\,,\quad\hat{\mu}_{T+1}(h)=\sqrt{\bar{\alpha}_{T}}\dot{\theta}_{T+1}\,,\quad\hat{\Sigma}_{T+1}(h)=\bar{\alpha}_{T}\dot{\Sigma}_{T+1}\,,$
where $\dot{\theta}_{T+1},\dot{\Sigma}_{T+1}\leftarrow{\tt IRLS}(\mathbf{0}_{d},I_{d}/\bar{\alpha}_{T},h)$. For $t\in[T]$, the distribution of $S_{t-1}\mid s_{t},h$ is
$p(s_{t-1}\mid s_{t},h)\approx\mathcal{N}(s_{t-1};\hat{\mu}_{t}(s_{t},h),\hat{\Sigma}_{t}(h))\,,\quad\hat{\mu}_{t}(s_{t},h)=\sqrt{\bar{\alpha}_{t-1}}\dot{\theta}_{t}\,,\quad\hat{\Sigma}_{t}(h)=\bar{\alpha}_{t-1}\dot{\Sigma}_{t}\,,$
where $\dot{\theta}_{t},\dot{\Sigma}_{t}\leftarrow{\tt IRLS}(\mu_{t}(s_{t})/\sqrt{\bar{\alpha}_{t-1}},\Sigma_{t}/\bar{\alpha}_{t-1},h)$.
Proof.
The proof is in Section A.4 and comprises four steps. First, we fix $t$ and apply approximation (6). Second, we reparameterize the prior, from a function of $s_{t}$ to a function of $s_{t}/\sqrt{\bar{\alpha}_{t}}$. Third, we combine the likelihood with the prior using the Laplace approximation. Finally, we repameterize the posterior, from a function of $s_{t}/\sqrt{\bar{\alpha}_{t}}$ to a function of $s_{t}$. ∎
Similarly to Theorem 2, the distributions in Theorem 4 mix evidence with the diffusion model prior. However, this is done implicitly in $\tt IRLS$. The posterior can be sampled from using $\tt LaplaceDPS$, where the mean and covariances would be taken from Theorem 4. Note that Theorem 2 is a special case of Theorem 4 where the mean function $g$ is an identity.
5 Application to Contextual Bandits
Algorithm 3 Contextual Thompson sampling.
1:for round $k=1,\dots,n$ do
2:     Sample $\tilde{\theta}_{k}\sim p(\cdot\mid h_{k})$, where $p(\cdot\mid h_{k})$ is the posterior distribution in (2)
3:     Take action $a_{k}\leftarrow\operatorname*{arg\,max\,}_{a\in\mathcal{A}}r(x_{k},a;\tilde{\theta}_{k})$ and observe reward $y_{k}$
Now we apply our posterior sampling approximations (Section 4) to contextual bandits. A contextual bandit [28, 31] is a classic model for sequential decision making under uncertainty where the agent takes actions conditioned on context. We denote the action set by $\mathcal{A}$ and the context set by $\mathcal{X}$. The mean reward of action $a\in\mathcal{A}$ in context $x\in\mathcal{X}$ is $r(x,a;\theta_{*})$, where $r:\mathcal{X}\times\mathcal{A}\times\Theta\to\mathbb{R}$ is some reward function and $\theta_{*}\in\Theta$ is an unknown model parameter (Section 2). The agent interacts with the bandit for $n$ rounds indexed by $k\in[n]$. In round $k$, the agent observes a context $x_{k}\in\mathcal{X}$, takes an action $a_{k}\in\mathcal{A}$, and observes a stochastic reward $y_{k}=r(x_{k},a_{k};\theta_{*})+\varepsilon_{k}$, where $\varepsilon_{k}$ is independent zero-mean $\sigma^{2}$-sub-Gaussian noise for $\sigma>0$. The goal of the agent is to maximize its cumulative reward in $n$ rounds, or equivalently to minimize its cumulative regret. We define the $n$-round regret as $R(n)=\sum_{k=1}^{n}\mathbb{E}\left[r(x_{k},a_{k,*};\theta_{*})-r(x_{k},a_{k};\theta_{*})\right]$, where $a_{k,*}$ is the optimal action in round $k$.
Arguably the most popular approach to solving contextual bandit problems is Thompson sampling [49, 11, 3]. The key idea in TS is to use the posterior distribution of $\theta_{*}$ to explore. This is done as follows. In round $k$, the model parameter is drawn from the posterior, $\tilde{\theta}_{k}\sim p(\cdot\mid h_{k})$ in (2), where $h_{k}$ is the history of all interactions up to round $k$. After that, the agent takes the action with the highest mean reward under $\tilde{\theta}_{k}$. The pseudo-code of this algorithm is given in Algorithm 3.
A linear bandit [13, 1] is a contextual bandit with a linear reward function $r(x,a;\theta_{*})=\phi(x,a)^{\top}\theta_{*}$, where $\phi:\mathcal{X}\times\mathcal{A}\to\mathbb{R}^{d}$ is a feature extractor. The feature extractor can be non-linear in $x$ and $a$. Therefore, linear bandits can be applied to non-linear functions in $x$ and $a$. The feature extractor can be hand-designed or learned [39]. To simplify notation, we let $\phi_{\ell}=\phi(x_{\ell},a_{\ell})$ be the feature vector of the action in round $k$. So the history of interactions up to round $k$ is $h_{k}=\left\{(\phi_{\ell},y_{\ell})\right\}_{\ell\in[k-1]}$. When the prior distribution is a Gaussian, $p(\theta_{*})=\mathcal{N}(\theta_{*};\theta_{0},\Sigma_{0})$, the posterior in round $k$ is a Gaussian in (3) for $h=h_{k}$. When the prior is a diffusion model, we propose sampling from
$\tilde{\theta}_{k}\leftarrow{\tt LaplaceDPS}((\mu_{t},\Sigma_{t})_{t\in[T]},h_{k})\,,$ (9)
where $\hat{\mu}_{t}$ and $\hat{\Sigma}_{t}$ in $\tt LaplaceDPS$ are set according to Theorem 2. We refer to the resulting algorithm as $\tt DiffTS$.
A generalized linear bandit [16, 23, 32, 25] is an extension of linear bandits to generalized linear models (Section 2.2). When $p(\theta_{*})=\mathcal{N}(\theta_{*};\theta_{0},\Sigma_{0})$, the Laplace approximation to the posterior is a Gaussian (Section 2.2). When the prior distribution is a diffusion model, we propose sampling from (9), where $\hat{\mu}_{t}$ and $\hat{\Sigma}_{t}$ in $\tt LaplaceDPS$ are set according to Theorem 4.
6 Experiments
Figure 2: Evaluation of $\tt DiffTS$ on three synthetic problems. The first row shows samples from the true (blue) and diffusion model (red) priors. The second row shows the regret of $\tt DiffTS$ and the baselines as a function of round $n$.
We conduct three experiments: synthetic problems in $2$ dimensions (Sections 6.2 and C.1), a recommender system (Section 6.3), and a classification problem (Section C.2). In addition, we conduct an ablation study in Section C.3, where we vary the number of training samples for the diffusion prior and the number of diffusion stages $T$.
6.1 Experimental Setup
We have four baselines. Three baselines are variants of contextual Thompson sampling [11, 3]: with an uninformative Gaussian prior ($\tt TS$), a learned Gaussian prior ($\tt TunedTS$), and a learned Gaussian mixture prior ($\tt MixTS$) [21]. The last baseline is diffusion posterior sampling ($\tt DPS$) of Chung et al. [12]. We implement all TS baselines as described in Section 5. The uninformative prior is $\mathcal{N}(\mathbf{0}_{d},I_{d})$. $\tt MixTS$ is used only in linear bandit experiments because the logistic regression variant does not exist. The TS baselines are chosen to cover various levels of prior information. Our implementation of $\tt DPS$ is described in Appendix D. We also experimented with frequentist baselines, such as $\tt LinUCB$ [1] and the $\varepsilon$-greedy policy. They performed worse than $\tt TS$ and thus we do not report them here.
Each experiment is set up as follows. First, the prior distribution of $\theta_{*}$ is specified: it can be synthetic or estimated from real-world data. Second, we learn this distribution from 10 0001000010\,00010 000 samples from it. In $\tt DiffTS$ and $\tt DPS$, we follow Appendix B. The number of stages is $T=100$ and the diffusion factor is $\alpha_{t}=0.97$. Since $0.97^{100}\approx 0.05$, most of the information in the training samples is diffused. The regressor in Appendix B is a $2$-layer neural network with ReLU activations. In $\tt TunedTS$, we fit the mean and covariance using maximum likelihood estimation. In $\tt MixTS$, we fit the Gaussian mixture using scikit-learn. All algorithms are evaluated on $\theta_{*}$ sampled from the true prior.
6.2 Synthetic Experiment
The first experiment is on three synthetic problems. Each problem is a linear bandit (Section 5) with $K=100$ actions in $d=2$ dimensions. The reward noise is $\sigma=2$. The feature vectors of actions are sampled uniformly at random from a unit ball. The prior distributions of $\theta_{*}$ are shown in Figure 2. The first is a mixture of two Gaussians and the last can be approximated well by a mixture of two Gaussians. $\tt MixTS$ is implemented with two mixture components, which can exactly represent the first prior and approximate the last one well.
Our results are reported in Figure 2. We observe two main trends. First, samples from the diffusion prior closely resemble those from the true prior. In such cases, $\tt DiffTS$ is expected to perform well and even outperforms $\tt MixTS$, because it has a better representation of the prior. We observe this in all problems. Second, $\tt DPS$ diverges as the number of rounds increases. This is because $\tt DPS$ relies on the score of the likelihood of observations (Section 7), which tends to infinity as the number of observations increases. This happens despite tuning (Appendix D). We report results on additional synthetic problems in Section C.1.
[…]
6.3 MovieLens Experiment
Figure 3: Evaluation of $\tt DiffTS$ on the MovieLens dataset.
In the second experiment, we learn to recommend an item to randomly arriving users. The problem is simulated using the MovieLens 1M dataset [27], with $1$ million ratings for 3 70637063\,7063 706 movies from 6 04060406\,0406 040 users. As a first step, we subtract the mean rating from all ratings and complete the sparse rating matrix $M$ by alternating least squares [14] with rank $d=5$. The learned factorization is $M=UV^{\top}$. The $i$-th row of $U$, denoted by $U_{i}$, represents user $i$. The $j$-th row of $V$, denoted by $V_{j}$, represents movie $j$. We use the item embeddings $V_{j}$ as unknown model parameters and the user embeddings $U_{i}$ as features of the actions.
We experiment with both linear and logistic bandits. In both, an item is initially chosen randomly from $V_{j}$ and $K=10$ random actions are chosen from $U_{i}$ in each round. In the linear bandit, the mean reward for recommending item $j$ to user $i$ is $U_{i}^{\top}V_{j}$. The reward noise is $\sigma=0.75$ and estimated from data. In the logistic bandit, the mean reward is $g(U_{i}^{\top}V_{j})$ where $g$ is a sigmoid.
Our MovieLens results are shown in Figure 3. We observe the same trends as in Section 6.2. First, samples from the diffusion prior closely resemble those from the true prior (Figure 3a). Since the problem is higher dimensional, we visualize the overlap using a UMAP projection [44]. Second, $\tt DiffTS$ has a significantly lower regret than all baselines, in both linear (Figure 3b) and logistic (Figure 3c) bandits. Finally, $\tt MixTS$ barely outperforms $\tt TunedTS$. We observe this trend consistently in higher dimensions, which motivated our work on online learning with more complex priors.
7 Related Work
There are two lines of related works: bandits with diffusion models and posterior sampling in diffusion models. We start with the former. Hsieh et al. [22] proposed Thompson sampling with a diffusion model prior for $K$-armed bandits. We consider a more general setting of contextual bandits and also develop new posterior sampling approximations for it. Aouali [4] proposed and analyzed contextual bandits with a linear diffusion model prior: $\mu_{t}(s_{t})$ in (5) is linear in $s_{t}$ and $q(s_{0})$ is a Gaussian. Due to this assumption, the resulting model is just a linear Gaussian model and not a general diffusion model, as in our work.
The closest related work on posterior sampling in diffusion models is $\tt DPS$ of Chung et al. [12]. The key idea in $\tt DPS$ is to sample from a diffusion model posterior by adding the score of the likelihood of observations to it. In our notation, this is $\nabla\log p(h\mid\theta)$, where $p(h\mid\theta)$ is defined in Assumptions 1 and 2. Note that $\nabla\log p(h\mid\theta)$ grows linearly in $N$. Therefore, $\tt DPS$ becomes unstable as $N\to\infty$. See our empirical results in Section 6.2 and the implementation of $\tt DPS$ in Appendix D, which was tuned to improve its stability. Other methods for posterior sampling in diffusion models have been proposed recently: a sequential Monte Carlo approximation for the conditional reverse process [52], a variant of $\tt DPS$ with an uninformative prior [37], a pseudo-inverse approximation to the likelihood of evidence [47], and posterior sampling in latent diffusion models [40]. All of these methods rely on the score of the likelihood $\nabla\log p(h\mid\theta)$ and thus become unstable as the number of observations $N$ increases. Our posterior approximations do not have this issue because they are based on the product of prior and evidence distributions (Theorems 2 and 4), and thus gradient-free. They work well across different levels of uncertainty (Section 6) and do not require tuning.
We also wanted to note that posterior sampling is a special from of guiding generation in diffusion models. Other approaches include conditional pre-training [15], a constraint in the reverse process [17], refining the null-space content [50], solving an optimization problem that pushes the reverse process towards evidence [46], and aligning the reverse process with the prompt [7].
8 Conclusions
We propose posterior sampling approximations for diffusion models priors. These approximations are contextual, and can be implemented efficiently in linear models and GLMs. We analyze them and evaluate them empirically on contextual bandit problems.
Our work focuses on developing novel posterior approximations rather than a regret analysis. This is because the existing approximations are unstable and may diverge in the online setting (Sections 7 and 6.2). A proper regret bound for $\tt DiffTS$ would require analyzing two errors. The first error arises because the reverse process reverses the forward process only approximately (Appendix B). The second error arises because our posteriors are only approximate (Section 4.3). That being said, prior works already provided general insights on the utility of complex priors. For instance, Russo and Van Roy [42] proved a $O(\sqrt{\Gamma H(A_{*})n})$ regret bound for a linear bandit, where $\Gamma$ is the maximum ratio of regret to information gain and $H(A_{*})$ is the entropy of the distribution of the optimal action under the prior. In short, the bound says that a lower entropy $H(A_{*})$, which reflects how concentrated the prior is, yields a lower regret.
Finally, we believe that our ideas can be extended beyond GLMs. The key idea in Section 4.4 is to use the Laplace approximation of the likelihood. This approximation can be computed exactly in GLMs. More generally though, it is a good approximation whenever the likelihood can be approximated well by a single Gaussian distribution. By the central limit theorem, under appropriate conditions, this is expected for any observation model when the number of observations is large.
Appendix A Proofs and Supplementary Lemmas
This section contains proofs of our main claims and supporting lemmas.
A.1 Proof of Lemma 1
All derivations are based on basic rules of probability and the chain structure in Figure 1, and are exact. From Figure 1, the joint probability distribution conditioned on $H=h$ factors as
$p(s_{0:T}\mid h)=p(s_{T}\mid h)\prod_{t=1}^{T}p(s_{t-1}\mid s_{t:T},h)=p(s_{T}\mid h)\prod_{t=1}^{T}p(s_{t-1}\mid s_{t},h)\,.$
We use that $p(s_{t-1}\mid s_{t:T},h)=p(s_{t-1}\mid s_{t},h)$ in the last equality. We consider two cases.
Derivation of $p(s_{t-1}\mid s_{t},h)$. By Bayes’ rule, we get
$p(s_{t-1}\mid s_{t},h)=\frac{p(h\mid s_{t-1},s_{t})\,p(s_{t-1}\mid s_{t})}{p(h\mid s_{t})}\propto p(h\mid s_{t-1})\,p(s_{t-1}\mid s_{t})\,.$
In the last step, we use that $p(h\mid s_{t})$ is a constant, since $s_{t}$ and $h$ are fixed, and that $p(h\mid s_{t-1},s_{t})=p(h\mid s_{t-1})$. Note that the last term $p(s_{t-1}\mid s_{t})$ is the conditional prior distribution. Let $t>1$. Then we rewrite the first term as
$p(h\mid s_{t-1})$ $=\int_{s_{0}}p(h,s_{0}\mid s_{t-1})\mathop{}\!\mathrm{d}s_{0}$
$=\int_{s_{0}}p(h\mid s_{0},s_{t-1})\,p(s_{0}\mid s_{t-1})\mathop{}\!\mathrm{d}s_{0}$
$=\int_{s_{0}}p(h\mid s_{0})\,p(s_{0}\mid s_{t-1})\mathop{}\!\mathrm{d}s_{0}\,.$
In the last equality, we use that our graphical model is a chain (Figure 1), and thus $p(h\mid s_{0},s_{t-1})=p(h\mid s_{0})$. Finally, we chain all identities and get that
$p(s_{t-1}\mid s_{t},h)\propto\int_{s_{0}}p(h\mid s_{0})\,p(s_{0}\mid s_{t-1})\mathop{}\!\mathrm{d}s_{0}\,p(s_{t-1}\mid s_{t})\,.$ (10)
Derivation of $p(s_{T}\mid h)$. By Bayes’ rule, we get
$p(s_{T}\mid h)=\frac{p(h\mid s_{T})\,p(s_{T})}{p(h)}\propto p(h\mid s_{T})\,p(s_{T})\,.$
In the last step, we use that $p(h)$ is a constant, since $h$ is fixed. The first term can be rewritten as
$p(h\mid s_{T})$ $=\int_{s_{0}}p(h,s_{0}\mid s_{T})\mathop{}\!\mathrm{d}s_{0}$
$=\int_{s_{0}}p(h\mid s_{0},s_{T})\,p(s_{0}\mid s_{T})\mathop{}\!\mathrm{d}s_{0}$
$=\int_{s_{0}}p(h\mid s_{0})\,p(s_{0}\mid s_{T})\mathop{}\!\mathrm{d}s_{0}\,.$
Finally, we chain all identities and get that
$p(s_{T}\mid h)\propto\int_{s_{0}}p(h\mid s_{0})\,p(s_{0}\mid s_{T})\mathop{}\!\mathrm{d}s_{0}\,p(s_{T})\,.$ (11)
This completes the derivations.
A.2 Proof of Theorem 2
This proof has two parts.
Derivation of $p(s_{t-1}\mid s_{t},h)$. From $p(h\mid s_{0})\propto\mathcal{N}(s_{0};\bar{\theta},\bar{\Sigma})$ and (6), it immediately follows that
$\int_{s_{0}}p(h\mid s_{0})\,p(s_{0}\mid s_{t-1})\mathop{}\!\mathrm{d}s_{0}\approx\mathcal{N}(s_{t-1}/\sqrt{\bar{\alpha}_{t-1}};\bar{\theta},\bar{\Sigma})\propto\mathcal{N}(s_{t-1};\sqrt{\bar{\alpha}_{t-1}}\bar{\theta},\bar{\alpha}_{t-1}\bar{\Sigma})\,.$
The last step treats $\bar{\alpha}_{t-1}$ and $\bar{\Sigma}$ as constants, because the forward process, $t$, and evidence are fixed. Now we apply Lemma 6 to distributions
$p(s_{t-1}\mid s_{t})=\mathcal{N}(s_{t-1};\mu_{t}(s_{t}),\Sigma_{t})\,,\quad\mathcal{N}(s_{t-1};\sqrt{\bar{\alpha}_{t-1}}\bar{\theta},\bar{\alpha}_{t-1}\bar{\Sigma})\,,$
and get that
$p(s_{t-1}\mid s_{t},h)\approx\mathcal{N}(s_{t-1};\hat{\mu}_{t}(s_{t},h),\hat{\Sigma}_{t}(h))\,,$
where $\hat{\mu}_{t}(s_{t},h)$ and $\hat{\Sigma}_{t}(h)$ are defined in the claim. This distribution is a product of two Gaussians: the prior with mean $\mu_{t}(s_{t})$ and covariance $\Sigma_{t}$, and the evidence with mean $\bar{\theta}$ and covariance $\bar{\Sigma}$.
Derivation of $p(s_{T}\mid h)$. Analogously to the derivation of $p(s_{t-1}\mid s_{t},h)$, we establish that
$\int_{s_{0}}p(h\mid s_{0})\,p(s_{0}\mid s_{T})\mathop{}\!\mathrm{d}s_{0}\approx\mathcal{N}(s_{T};\sqrt{\bar{\alpha}_{T}}\bar{\theta},\bar{\alpha}_{T}\bar{\Sigma})\,.$
Then we apply Lemma 6 to distributions
$p(s_{T})=\mathcal{N}(s_{T};\mathbf{0}_{d},I_{d})\,,\quad\mathcal{N}(s_{T};\sqrt{\bar{\alpha}_{T}}\bar{\theta},\bar{\alpha}_{T}\bar{\Sigma})\,,$
and get that
$p(s_{T}\mid h)\approx\mathcal{N}(s_{T};\hat{\mu}_{T+1}(h),\hat{\Sigma}_{T+1}(h))\,,$
where $\hat{\mu}_{T+1}(h)$ and $\hat{\Sigma}_{T+1}(h)$ are given in the claim. This distribution is a product of two Gaussians: the prior with mean $\mathbf{0}_{d}$ and covariance $I_{d}$, and the evidence with mean $\bar{\theta}$ and covariance $\bar{\Sigma}$.
A.3 Proof of Theorem 3
We start with the triangle inequality
$\|\tilde{\theta}-\theta_{*}\|_{2}=\|\tilde{\theta}-\bar{\theta}+\bar{\theta}-\theta_{*}\|_{2}\leq\|\tilde{\theta}-\bar{\theta}\|_{2}+\|\bar{\theta}-\theta_{*}\|_{2}\,,$
where we introduce $\bar{\theta}$ from Section 2.1. Now we bound each term on the right-hand side.
Upper bound on $\|\tilde{\theta}-\bar{\theta}\|_{2}$. This part of the proof is based on analyzing the asymptotic behavior of the conditional densities in Theorem 2.
As a first step, note that $S_{T}\sim\mathcal{N}(\hat{\mu}_{T+1}(h),\hat{\Sigma}_{T+1}(h))$, where
[…]
Since $\lambda_{d}(\bar{\Sigma}^{-1})\to\infty$, we get
$\hat{\Sigma}_{T+1}(h)\to\bar{\alpha}_{T}\bar{\Sigma}\,,\quad\hat{\mu}_{T+1}(h)\to\sqrt{\bar{\alpha}_{T}}\bar{\theta}\,.$
Moreover, $\lambda_{d}(\bar{\Sigma}^{-1})\to\infty$ implies $\lambda_{1}(\bar{\Sigma})\to 0$, and thus $\lim_{N\to\infty}\|S_{T}-\sqrt{\bar{\alpha}_{T}}\bar{\theta}\|_{2}=0$.
The same argument can be applied inductively to later stages of the reverse process. Specifically, for any $t\in[T]$, $S_{t-1}\sim\mathcal{N}(\hat{\mu}_{t}(S_{t},h),\hat{\Sigma}_{t}(h))$, where
$\hat{\mu}_{t}(S_{t},h)=\hat{\Sigma}_{t}(h)(\Sigma_{t}^{-1}\mu_{t}(S_{t})+\bar{\Sigma}^{-1}\bar{\theta}/\sqrt{\bar{\alpha}_{t-1}})\,,\quad\hat{\Sigma}_{t}(h)=(\Sigma_{t}^{-1}+\bar{\Sigma}^{-1}/\bar{\alpha}_{t-1})^{-1}\,.$
Since $\lambda_{d}(\bar{\Sigma}^{-1})\to\infty$ and $S_{t}\to\sqrt{\bar{\alpha}_{t}}\bar{\theta}$ by induction, we get
$\hat{\Sigma}_{t}(h)\to\bar{\alpha}_{t-1}\bar{\Sigma}\,,\quad\hat{\mu}_{t}(S_{t},h)\to\sqrt{\bar{\alpha}_{t-1}}\bar{\theta}\,.$
Moreover, $\lambda_{d}(\bar{\Sigma}^{-1})\to\infty$ implies $\lambda_{1}(\bar{\Sigma})\to 0$, and thus $\lim_{N\to\infty}\|S_{t-1}-\sqrt{\bar{\alpha}_{t-1}}\bar{\theta}\|_{2}=0$.
In the last stage, $t=1$, $\bar{\alpha}_{0}=1$, and $S_{0}=\tilde{\theta}$, which implies that
$\lim_{N\to\infty}\|\tilde{\theta}-\bar{\theta}\|_{2}\to 0\,.$
Upper bound on $\|\bar{\theta}-\theta_{*}\|_{2}$. This part of the proof uses the definition of $\bar{\theta}$ in Section 2.1 and that $\varepsilon_{\ell}\sim\mathcal{N}(0,\sigma^{2})$ is independent noise. By definition,
$\bar{\theta}-\theta_{*}=\sigma^{-2}\bar{\Sigma}\sum_{\ell=1}^{N}\phi_{\ell}y_{\ell}-\theta_{*}=\sigma^{-2}\bar{\Sigma}\sum_{\ell=1}^{N}\phi_{\ell}(\phi_{\ell}^{\top}\theta_{*}+\varepsilon_{\ell})-\theta_{*}=\sigma^{-2}\bar{\Sigma}\sum_{\ell=1}^{N}\phi_{\ell}\varepsilon_{\ell}\,.$
Since $\varepsilon_{\ell}$ is independent zero-mean Gaussian noise with variance $\sigma^{2}$, $\bar{\theta}-\theta_{*}$ is a Gaussian random variable with mean $\mathbf{0}_{d}$ and covariance
$\mathrm{cov}\left[\sigma^{-2}\bar{\Sigma}\sum_{\ell=1}^{N}\phi_{\ell}\varepsilon_{\ell}\right]=\sigma^{-4}\bar{\Sigma}\left(\sum_{\ell=1}^{N}\phi_{\ell}\mathrm{var}\left[\varepsilon_{\ell}\right]\phi_{\ell}^{\top}\right)\bar{\Sigma}=\bar{\Sigma}\frac{\sum_{\ell=1}^{N}\phi_{\ell}\phi_{\ell}^{\top}}{\sigma^{2}}\bar{\Sigma}=\bar{\Sigma}\,.$
Since $\lambda_{d}(\bar{\Sigma}^{-1})\to\infty$ implies $\lambda_{1}(\bar{\Sigma})\to 0$, we get
$\lim_{N\to\infty}\|\bar{\theta}-\theta_{*}\|_{2}=0\,.$
This completes the proof.
A.4 Proof of Theorem 4
This proof has two parts.
Derivation of $p(s_{t-1}\mid s_{t},h)$. From (6), we have
$\int_{s_{0}}p(h\mid s_{0})\,p(s_{0}\mid s_{t-1})\mathop{}\!\mathrm{d}s_{0}\approx p(h\mid s_{t-1}/\sqrt{\bar{\alpha}_{t-1}})\,.$
Since $p(s_{t-1}\mid s_{t})$ is a Gaussian, we have
$p(s_{t-1}\mid s_{t})=\mathcal{N}(s_{t-1};\mu_{t}(s_{t}),\Sigma_{t})\propto\mathcal{N}(\gamma s_{t-1};\gamma\mu_{t}(s_{t}),\gamma^{2}\Sigma_{t})$
for $\gamma=1/\sqrt{\bar{\alpha}_{t-1}}$. Then by the Laplace approximation,
$p(h\mid\gamma s_{t-1})\,\mathcal{N}(\gamma s_{t-1};\gamma\mu_{t}(s_{t}),\gamma^{2}\Sigma_{t})=\mathcal{N}(\gamma s_{t-1};\dot{\theta}_{t},\dot{\Sigma}_{t})\propto\mathcal{N}(s_{t-1};\dot{\theta}_{t}/\gamma,\dot{\Sigma}_{t}/\gamma^{2})\,,$
where $\dot{\theta}_{t},\dot{\Sigma}_{t}\leftarrow{\tt IRLS}(\gamma\mu_{t}(s_{t}),\gamma^{2}\Sigma_{t},h)$.
Derivation of $p(s_{T}\mid h)$. Analogously to the derivation of $p(s_{t-1}\mid s_{t},h)$, we establish that
$\int_{s_{0}}p(h\mid s_{0})\,p(s_{0}\mid s_{T})\mathop{}\!\mathrm{d}s_{0}\approx p(h\mid s_{T}/\sqrt{\bar{\alpha}_{T}})\,.$
Then by the Laplace approximation for $\gamma=1/\sqrt{\bar{\alpha}_{T}}$, we get
$p(h\mid\gamma s_{t-1})\,\mathcal{N}(s_{t-1};\mathbf{0}_{d},I_{d})\propto\mathcal{N}(s_{t-1};\dot{\theta}_{T+1}/\gamma,\dot{\Sigma}_{T+1}/\gamma^{2})$
where $\dot{\theta}_{T+1},\dot{\Sigma}_{T+1}\leftarrow{\tt IRLS}(\mathbf{0}_{d},\gamma^{2}I_{d},h)$.
A.5 Supplementary Lemmas
We state and prove our supplementary lemmas next.
Lemma 5.
Let $p(x)=\mathcal{N}(x;\mu_{1},\Sigma_{1})$ and $q(x)=\mathcal{N}(x;\mu_{2},\Sigma_{2})$, where $\mu_{1},\mu_{2}\in\mathbb{R}^{d}$ and $\Sigma_{1},\Sigma_{2}\in\mathbb{R}^{d\times d}$. Then
$d(p,q)=\frac{1}{2}\left((\mu_{2}-\mu_{1})^{\top}\Sigma_{2}^{-1}(\mu_{2}-\mu_{1})+\operatorname{tr}(\Sigma_{2}^{-1}\Sigma_{1})-\log\frac{\operatorname{det}(\Sigma_{1})}{\operatorname{det}(\Sigma_{2})}-d\right)\,.$
Moreover, when $\Sigma_{1}=\Sigma_{2}$,
$d(p,q)=\frac{1}{2}(\mu_{2}-\mu_{1})^{\top}\Sigma_{2}^{-1}(\mu_{2}-\mu_{1})\,.$
Proof.
The proof follows from the definitions of KL divergence and multivariate Gaussians. ∎
Lemma 6.
Fix $\mu_{1}\in\mathbb{R}^{d}$, $\Sigma_{1}\succeq 0$, $\mu_{2}\in\mathbb{R}^{d}$, and $\Sigma_{2}\succeq 0$. Then
$\mathcal{N}(x;\mu_{1},\Sigma_{1})\,\mathcal{N}(x;\mu_{2},\Sigma_{2})\propto\mathcal{N}(x;\mu,\Sigma)\,,$
where
$\mu=\Sigma(\Sigma_{1}^{-1}\mu_{1}+\Sigma_{2}^{-1}\mu_{2})\,,\quad\Sigma=(\Sigma_{1}^{-1}+\Sigma_{2}^{-1})^{-1}\,.$
Proof.
This is a classic result, which is proved as
$\mathcal{N}(x;\mu_{1},\Sigma_{1})\,\mathcal{N}(x;\mu_{2},\Sigma_{2})$ $\propto\exp\left[-\frac{1}{2}((x-\mu_{1})^{\top}\Sigma_{1}^{-1}(x-\mu_{1})+(x-\mu_{2})^{\top}\Sigma_{2}^{-1}(x-\mu_{2}))\right]$
$\propto\exp\left[-\frac{1}{2}(x^{\top}\Sigma_{1}^{-1}x-2x^{\top}\Sigma_{1}^{-1}\mu_{1}+x^{\top}\Sigma_{2}^{-1}x-2x^{\top}\Sigma_{2}^{-1}\mu_{2})\right]$
$=\exp\left[-\frac{1}{2}(x^{\top}\Sigma^{-1}x-2x^{\top}\Sigma^{-1}\Sigma(\Sigma_{1}^{-1}\mu_{1}+\Sigma_{2}^{-1}\mu_{2}))\right]$
$\propto\exp\left[-\frac{1}{2}(x-\mu)^{\top}\Sigma^{-1}(x-\mu)\right]\propto\mathcal{N}(x;\mu,\Sigma)\,.$
The neglected factors depend on $\mu_{1}$, $\mu_{2}$, $\Sigma_{1}$, and $\Sigma_{2}$. This completes the proof. ∎
Appendix B Learning the Reverse Process
One property of our model is that $q(s_{T})\approx\mathcal{N}(s_{T};\mathbf{0}_{d},I_{d})$ when $T$ is sufficiently large [18]. Since $S_{T}$ is initialized to the same distribution in the reverse process $p$, $p$ can be learned from the forward process $q$ by simply reversing it. This is done as follows. Using the definition of the forward process in (4), Ho et al. [18] showed that
$q(s_{t-1}\mid s_{t},s_{0})=\mathcal{N}(s_{t-1};\tilde{\mu}_{t}(s_{t},s_{0}),\tilde{\beta}_{t}I_{d})$ (12)
holds for any $s_{0}$ and $s_{t}$, where
$\tilde{\mu}_{t}(s_{t},s_{0})=\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}s_{0}+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}s_{t}\,,\quad\tilde{\beta}_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}\,,\quad\bar{\alpha}_{t}=\prod_{\ell=1}^{t}\alpha_{\ell}\,.$ (13)
Therefore, the latent variable in stage $t-1$, $S_{t-1}$, is easy to sample when $s_{t}$ and $s_{0}$ are known. To estimate $s_{0}$, which is unknown when sampling from the reverse process, we use the forward process again. In particular, (4) implies that $s_{t}=\sqrt{\bar{\alpha}_{t}}s_{0}+\sqrt{1-\bar{\alpha}_{t}}\tilde{\varepsilon}_{t}$ for any $s_{0}$, where $\tilde{\varepsilon}_{t}\sim\mathcal{N}(\mathbf{0}_{d},I_{d})$ is a standard Gaussian noise. This identity can be rearranged as
$s_{0}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}(s_{t}-\sqrt{1-\bar{\alpha}_{t}}\tilde{\varepsilon}_{t})\,,$
To obtain $\tilde{\varepsilon}_{t}$, which is unknown when sampling from $p$, we learn to regress it from $s_{t}$ [18].
The regressor is learned as follows. Let $\varepsilon_{t}(\cdot;\psi)$ be a regressor of $\tilde{\varepsilon}_{t}$ parameterized by $\psi$ and $\mathcal{D}=\left\{s_{0}\right\}$ be a dataset of training examples. We sample $s_{0}$ uniformly at random from $\mathcal{D}$ and then solve
$\psi_{t}=\operatorname*{arg\,min\,}_{\psi}\mathbb{E}_{q}\left[\|\tilde{\varepsilon}_{t}-\varepsilon_{t}(S_{t};\psi)\|_{2}^{2}\right]$ (14)
per stage. The expectation is approximated by an empirical average over sampled $s_{0}$. Note that we slightly depart from Ho et al. [18]. Since each $\varepsilon_{t}$ has its own parameters, the original optimization problem over $T$ stages decomposes into $T$ subproblems.
Appendix C Additional Experiments
This section contains three additional experiments.
C.1 Additional Synthetic Problems
Figure 4: Evaluation of $\tt DiffTS$ on another three synthetic problems. The first row shows samples from the true (blue) and diffusion model (red) priors. The second row shows the regret of $\tt DiffTS$ and the baselines as a function of round $n$.
In Section 6.2, we show results for three hand-selected problems out of six. We report results on the other three problems in Figure 4. We observe the same trends as in Section 6.2. The only minor change is that $\tt DPS$ does not diverge on the last problem.
C.2 MNIST Experiment
Figure 5: Evaluation of $\tt DiffTS$ on the MNIST dataset.
The next experiment is on the MNIST dataset [30], where the mean reward is estimated by a classifier. We start with learning an MLP-based multi-way classifier for digits and extract $d=8$ dimensional embeddings of all digits in the dataset, which are used as feature vectors in our experiment. We generate a distribution over model parameters $\theta_{*}$ as follows: (1) we choose a random positive label, assign it reward $1$, and assign reward $-1$ to all other labels; (2) we subsample a random dataset of size $20$, with $50\%$ positive and $50\%$ negative labels; (3) we train a linear model, which gives us a single $\theta_{*}$. We repeat this 10 0001000010\,00010 000 times and get a distribution over $\theta_{*}$.
We consider both linear and logistic bandits. In both, the model parameter $\theta_{*}$ is initially sampled from the prior. In each round, $K=10$ random actions are chosen randomly from all digits. In the linear bandit, the mean reward for digit $x$ is $x^{\top}\theta_{*}$ and the reward noise is $\sigma=1$. In the logistic bandit, the mean reward is $g(x^{\top}\theta_{*})$ where $g$ is a sigmoid.
Our MNIST results are reported in Figure 5. We observe again that $\tt DiffTS$ has a lower regret than all baselines, because the learned prior captures the underlying distribution of $\theta_{*}$ well. We note that both the prior and diffusion prior distributions exhibit a strong cluster structure (Figure 5a), where each cluster represents one label.
C.3 Ablation Studies
Figure 6: An ablation study of $\tt DiffTS$ on the cross problem: (a) regret with a varying number of samples for training the diffusion prior, (b) regret with a varying number of diffusion stages $T$, and (c) computation time with a varying number of diffusion stages $T$.
We conduct three ablation studies on the cross problem in Figure 2.
In all experiments, the number of samples for training diffusion priors was 10 0001000010\,00010 000. In Figure 6a, we vary it from $100$ to 10 0001000010\,00010 000. We observe that the regret decreases as the number of samples increases, due to learning a more accurate prior approximation. It stabilizes around 3 00030003\,0003 000 training samples. We conclude that the quality of the learned prior approximation has a major impact on regret.
In all experiments, the number of diffusion stages was $T=100$. In Figure 6b, we vary it from $1$ to $300$ and observe its impact on regret. The regret at $T=1$ is high, but then it drops quickly up to $T=25$. It stabilizes around $T=100$, which we used in our experiments. In Figure 6c, we vary $T$ from $1$ to $300$ and observe its impact on the computation time of posterior sampling. As suggested in Section 4.2, it increases linearly with $T$. The main contribution to the computation time are $T$ calls to the neural network regressor.
Appendix D Implementation of Chung et al. [12]
Algorithm 4 $\tt DPS$ of Chung et al. [12].
1:Input: Model parameters $\tilde{\sigma}_{t}$ and $\zeta_{t}$
2:
3:Initial sample $S_{T}\sim\mathcal{N}(\mathbf{0}_{d},I_{d})$
4:for stage $t=T,\dots,1$ do
5:     $\hat{S}\leftarrow\frac{\varepsilon_{t}(S_{t};\psi_{t})}{\sqrt{1-\bar{\alpha}_{t}}}$
6:     $\hat{S}_{0}\leftarrow\frac{1}{\sqrt{\bar{\alpha}_{t}}}(S_{t}+(1-\bar{\alpha}_{t})\hat{S})$
7:     $Z\sim\mathcal{N}(\mathbf{0}_{d},I_{d})$
8:     $S_{t-1}\leftarrow\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\hat{S}_{0}+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}S_{t}+\tilde{\sigma}_{t}Z-\zeta_{t}\nabla\sum_{\ell=1}^{N}(y_{\ell}-\phi_{\ell}^{\top}\hat{S}_{0})^{2}$
9:
10:Output: Posterior sample $S_{0}$
In our experiments, we compare to diffusion posterior sampling ($\tt DPS$) with a Gaussian observation noise (Algorithm 1 in Chung et al. [12]). We implement it as shown in Algorithm 4. The score is $\hat{S}=\varepsilon_{t}(S_{t};\psi_{t})/\sqrt{1-\bar{\alpha}_{t}}$ because we regress $\tilde{\varepsilon}_{t}$ in Appendix B. We set $\tilde{\sigma}_{t}=\sqrt{\tilde{\beta}_{t}}$, which is the same noise as in our reverse process (Section 3). The term
$\nabla\sum_{\ell=1}^{N}(y_{\ell}-\phi_{\ell}^{\top}\hat{S}_{0})^{2}$
is the gradient of the negative log-likelihood with respect to $S_{t}$ evaluated at $\hat{S}_{0}$.
As noted in Appendices C.2 and D.1 of Chung et al. [12], $\zeta_{t}$ in $\tt DPS$ needs to be tuned for good performance. We also observed this in our experiments (Section 6.2). To make $\tt DPS$ work well, we follow Chung et al. [12] and set
$\zeta_{t}=\frac{1}{\sqrt{\sum_{\ell=1}^{N}(y_{\ell}-\phi_{\ell}^{\top}\hat{S}_{0})^{2}}}\,.$
While this significantly improves the performance of $\tt DPS$, it does not prevent failures. The fundamental problem is that gradient-based optimization is sensitive to the step size, especially when the optimized function is steep. We do not have any such hyper-parameter in our algorithm.